OpenStudy (anonymous):

Now its addition logx+log7=39

OpenStudy (anonymous):

is it really 39 cross check it it must be 49

OpenStudy (anonymous):

Nope it says 39.

OpenStudy (anonymous):

rule is logx +logy means log (x * y)

OpenStudy (anonymous):

so 7x=39?

OpenStudy (turingtest):

logx+log7=39 log7x=39 7x=10^39 x=1.43X10^38

OpenStudy (anonymous):

turning test don't give answer straight a way first explain the rules it will help him to understand the things in abetter way

OpenStudy (anonymous):

Yeah im lost by the x=1.43X10^38

OpenStudy (anonymous):

yeah but here log7x=39

OpenStudy (turingtest):

you already gave the rules sheg, otherwise I would have My anser is in scientific notation since the number is so large (assuming a base 10 log) log7x=39 7x=10^39 x=(10^39)/7=1.43(10^38) I just wrote the answer inscientic notation because the number is so large

OpenStudy (anonymous):

the answer would be like. 5.55

OpenStudy (anonymous):

now if we will take antilog we will get 7x = antilog (39)

OpenStudy (anonymous):

I dont know what antilog is ?

OpenStudy (turingtest):

neither do I LOL!

OpenStudy (anonymous):

it is inverse of log!!

OpenStudy (turingtest):

well I figure as much...

OpenStudy (turingtest):

just never heard of it before and don't see why you need it

OpenStudy (anonymous):

because u have log only on one side not on both the sides

OpenStudy (anonymous):

Alright. Thank you again! :)

OpenStudy (anonymous):

hey friend here if ur answer is 5.5 then ur question must be logx + log 7 = log 39